### v1

# On the number of permutation-twisted dot products

Abstract.
For distinct real numbers $a_{1},\ldots,a_{n}$ and distinct real numbers $b_{1},\ldots,b_{n}$, consider the sum $S=\sum_{i=1}^{n}a_{i}b_{\pi(i)}$ as $\pi$ ranges over the permutations of $[n]$. We show that this sum always assumes at least $\Omega(n^{3})$ distinct values, which is optimal. This “support” bound complements recent work of Do, Nguyen, Phan, Tran, and Vu on the anticoncentration properties of $S$ when $\pi$ is chosen uniformly at random.
1. Introduction
Let $\mathbf{a}=(a_{1},\ldots,a_{n})$ and $\mathbf{b}=(b_{1},\ldots,b_{n})$ be $n$-tuples of real numbers. For each permutation $\pi$ in the symmetric group $\mathfrak{S}_{n}$, define the quantity
$S(\mathbf{a},\mathbf{b};\pi):=\sum_{i=1}^{n}a_{i}b_{\pi(i)},$
which is the “dot product of $\mathbf{a},\mathbf{b}$ twisted by $\pi$”. Motivated by Littlewood–Offord theory and a geometric problem of Pawlowski [P24], Do, Nguyen, Phan, Tran, and Vu [DNPTV25] recently raised the problem of studying the anticoncentration properties of the random variable $S(\mathbf{a},\mathbf{b};\pi)$ where $\pi$ is chosen uniformly at random. One of their main results states that if the $a_{i}$’s are all distinct and the $b_{i}$’s are all distinct, then each possible value of the sum is assumed with probability at most $O(n^{-5/2}\log n)$; this bound is optimal up to the logarithm (which likely can be removed). This work has prompted a small flurry of activity [ABDV26, HPZ26].
We are interested in the set
$\mathcal{S}(\mathbf{a},\mathbf{b}):=\left\{S(\mathbf{a},\mathbf{b};\pi):\pi\in\mathfrak{S}_{n}\right\}$
of all possible sums as $\pi$ ranges; this is of course precisely the support of Do, Nguyen, Phan, Tran, and Vu’s random variable $S(\mathbf{a},\mathbf{b};\pi)$. Our main result is that if the $a_{i}$’s are all distinct and the $b_{i}$’s are all distinct, then there are at least $\Omega(n^{3})$ distinct sums.
Theorem 1.1.
There exists an absolute constant $c>0$ such that $|\mathcal{S}(\mathbf{a},\mathbf{b})|\geq cn^{3}$ for all $\mathbf{a},\mathbf{b}\in\mathbb{R}^{n}$ with distinct entries.
Some version of the distinctness condition on the $a_{i}$’s and $b_{i}$’s is necessary since otherwise $S(\mathbf{a},\mathbf{b};\pi)$ may not vary at all as $\pi$ varies. The bound in Theorem 1.1 is sharp up to the value of the implicit constant $c$ since $|\mathcal{S}(\mathbf{a},\mathbf{b})|=(1/6+o(1))n^{3}$ for $\mathbf{a}=\mathbf{b}=(1,2,\ldots,n)$. We remark that one can obtain the trivial lower bound $|\mathcal{S}(\mathbf{a},\mathbf{b})|\geq 1+\binom{n}{2}$ by first ordering both tuples $\mathbf{a},\mathbf{b}$ in increasing order and then iteratively swapping a pair of adjacent elements of $\mathbf{b}$ until all elements are in decreasing order.
Our proof of Theorem 1.1 is elementary and fairly short. As it turns out, we are able to obtain $\Omega(n^{3})$ distinct sums using a somewhat small set of permutations, namely, the permutations obtained as an involution times a fixed “base” permutation. Pohoata [P26] has recently obtained an independent proof of Theorem 1.1 using more complicated methods.
2. Proofs
Let $\mathbf{a},\mathbf{b}\in\mathbb{R}^{n}$ be $n$-tuples each with distinct entries. Write $S_{0}:=S(\mathbf{a},\mathbf{b};\operatorname{id})$, where $\operatorname{id}$ denotes the identity permutation. For the permutation $\pi=(j,k)$ that swaps $j,k$ and fixes all elements of $n]\setminus\{j,k\}$, we can compute
$S(\mathbf{a},\mathbf{b};(j,k))=S_{0}+a_{j}(b_{k}-b_{j})+a_{k}(b_{j}-b_{k})=S_{0}-(a_{k}-a_{j})(b_{k}-b_{j}).$
In general, if $\pi=(j_{1},k_{1})\cdots(j_{r},k_{r})$ is a product of disjoint transpositions, then we obtain
$S(\mathbf{a},\mathbf{b};(j_{1},k_{1})\cdots(j_{r},k_{r}))=S_{0}-\sum_{i=1}^{r}(a_{k_{i}}-a_{j_{i}})(b_{k_{i}}-b_{j_{i}}).$
Thus, to prove Theorem 1.1, it suffices to find disjoint transpositions $(j_{1},k_{1}),\ldots,(j_{r},k_{r})$ with the property that the set
(1) $\{(a_{k_{1}}-a_{j_{1}})(b_{k_{1}}-b_{j_{1}}),\ldots,(a_{k_{r}}-a_{j_{r}})(b_{k_{r}}-b_{j_{r}})\}$
has $\Omega(n^{3})$ distinct subset sums. We are of course free to reorder the entries of $\mathbf{a}$ and of $\mathbf{b}$ (possibly in different ways) before specifying our family of disjoint transpositions.
In general, a set of $m$ real numbers can have as few as $O(m^{2})$ distinct subset sums (one example is $\{1,\dots,m\}$). To get a cubic lower bound for the set (1), we will choose transpositions that capture the behavior of $\mathbf{a}$ and $\mathbf{b}$ at many different scales.
2.1. Constructing the involutions
The first step is identifying a family of disjoint pairs $(j_{i},k_{i})$ of elements of $[n]$ such that the corresponding differences $a_{k_{i}}-a_{j_{i}}$ “see” the various scales of $\mathbf{a}$; and likewise for $\mathbf{b}$. We achieve this iteratively, as follows.
Without loss of generality, we may assume that $a_{1}<\cdots<a_{n}$. To begin, among all choices of $1\leq j<k\leq n$, choose a pair minimizing the quantity $a_{k}-a_{j}$, and let $(j_{1},k_{1})$ denote this pair. Suppose we have already obtained disjoint pairs $(j_{1},k_{1}),\ldots,(j_{s},k_{s})$ (with each $j_{i}<k_{i}$). Let $\mathcal{U}_{s}$ be the set of n2sn-2s unpaired indices of $[n]$. Among all choices of $j<k$ in $\mathcal{U}_{s}$ with
$|\{j+1,\dots,k-1\}\cap\mathcal{U}_{s}|=s-1,$
choose one that minimizes $a_{k}-a_{j}$, and let $(j_{s+1},k_{s+1})$ denote this pair; if there are no such pairs, then halt the procedure.
Suppose that the procedure produced the pairs $(j_{1},k_{1}),\ldots,(j_{r},k_{r})$ and then halted. Since it failed to produce $(j_{r+1},k_{r+1})$, we must have had $n-2r=|\mathcal{U}_{r}|\leq r+1$, i.e., $r\geq(n-1)/3$. Thus, our process always produces a linear number of pairs. For each $1\leq i\leq r$, write $x(i):=a_{k_{i}}-a_{j_{i}}$. The crucial property of the sequence $x(1),\ldots,x(r)$ (besides its length) is its superadditivity.
Lemma 2.1 (Superadditivity).
We have $x(u+v)\geq x(u)+x(v)$ for all $1\leq u,v\leq r$ with $u+v\leq r$.
Proof.
Write $i_{1}<i_{2}<\cdots<i_{u+v-1}$ for the elements of $\mathcal{U}_{u+v}$ between $j_{u+v}$ and $k_{u+v}$. All of these elements are also contained in $\mathcal{U}_{u}$ and $\mathcal{U}_{v}$, so the greedy nature of our inductive procedure guarantees that $x(u)\leq a_{i_{u}}-a_{j_{u+v}}$ and $x(v)\leq a_{k_{u+v}}-a_{i_{u}}$, whence $x(u)+x(v)\leq a_{k_{u+v}}-a_{j_{u+v}}=x(u+v)$. ∎
In the same way we obtain a superadditive sequence $y(1),\ldots,y(r^{\prime})$ from $\mathbf{b}$, again with $r^{\prime}=\Omega(n)$. It remains to show that the set of numbers $s(i):=x(i)y(i)$, for $i$ between $1$ and $\min(r,r^{\prime})=\Theta(n)$, has $\Omega(n^{3})$ distinct subset sums.
2.2. Subset sums for dot products of superadditive sequences
For each $m\geq 1$, let $f(m)$ denote the minimum possible number of distinct subset sums of $\{s(1),\ldots,s(m)\}$, where $x(1),\ldots,x(m)$ and $y(1),\dots,y(m)$ are superadditive sequences of positive real numbers and $s(i):=x(i)y(i)$. In light of the preceding discussion, the reader will be unsurprised to see that our main technical lemma takes the following form.
Lemma 2.2.
There is a constant $\kappa>0$ such that $f(m)\geq\kappa m^{3}$ for all $m\geq 1$.
Proof.
Let $x(1),\ldots,x(m)$ and $y(1),\ldots,y(m)$ be superadditive sequences of positive reals, and set $s(i):=x(i)y(i)$. The key claim is that there exists $t(m)=(1-O(m^{-1/3}))\cdot(3m^{2})^{1/3}$ such that
(2) $\sum_{i=1}^{t(m)}s(i)<s(m).$
Before proving this claim, let us see how it implies the conclusion of the lemma.
All subset sums of $\{s(1),\dots,s(m-1)\}$ are also subset sums of $\{s(1),\dots,s(m)\}$. Moreover, for each $S\subseteq\{1,\dots,t(m)\}$, the subset sum
$\sum_{i\in\{1,\dots,m\}\setminus S}s(i)=\sum_{i=1}^{m-1}s(i)-\sum_{j\in S}s(j)+s(m)$
is larger than $s_{1}+\cdots+s_{m-1}$, so it is a subset sum of $\{s(1),\ldots,s(m)\}$ that is not a subset sum of $\{s(1),\ldots,s(m-1)\}$. By varying $S$, we obtain at least $f(t(m))$ new subset sums, so
(3) $f(m)\geq f(m-1)+f(t(m)).$
Let $\kappa_{m}:=m^{-3}f(m)$. We show that $\kappa_{m}$ is uniformly bounded away from 0. By the recurrence (3),
$f(m)$ $\geq f(m-1)+f(t(m))$
$\geq\kappa_{m-1}(m-1)^{3}+\kappa_{t(m)}\left((1-O(m^{-1/3}))\cdot(3m^{2})^{1/3}\right)^{3}$
$\geq\min\left\{\kappa_{m-1},(1-O(m^{-1/3}))\cdot\kappa_{t(m)}\right\}\cdot((m-1)^{3}+3m^{2})$
$\geq\min\left\{\kappa_{m-1},\exp\left(-O(m^{-1/3})\right)\cdot\kappa_{t(m)}\right\}\cdot m^{3},$
so
$\kappa_{m}\geq\min\left\{\kappa_{m-1},\exp\left(-O(m^{-1/3})\right)\cdot\kappa_{t(m)}\right\}.$
Thus, even if $\kappa_{1},\kappa_{2},\kappa_{3},\dots$ does decrease, it does so extremely slowly. Concretely, there is some natural number $m^{*}$ such that for any $m$, the sequence $m_{0}:=m,~m_{1}:=t(m_{0}),~m_{2}:=t(m_{1}),$$\ldots$ eventually reaches some $m_{q}\leq m^{*}$. In this case we can bound
$\kappa_{m}\geq\prod_{i=0}^{q-1}\exp\left(-O(m_{i}^{-1/3})\right)\cdot\min_{1\leq m^{\prime}\leq m^{*}}\kappa_{m^{\prime}}\gg\exp\left(-\sum_{i=0}^{q-1}O(m_{i}^{-1/3})\right).$
As $i$ decreases, the sequence $m_{i}$ grows doubly-exponentially, so the product on the right-hand side is uniformly bounded below by some $\kappa>0$.
It remains to prove the main claim. The intuition is that for $m=\left(\sum_{i=1}^{t}i^{2}\right)^{1/2}\approx(t^{3}/3)^{1/2}$, two “fractional” applications of superadditivity morally give the chain of inequalities
$\sum_{i=1}^{t}x(i)y(i)\leq\sum_{i=1}^{t}\frac{i}{m}x(m)\cdot y(i)\leq x(m)y\left(\sum_{i=1}^{t}\frac{i^{2}}{m}\right)=x(m)y(m).$
To make this heuristic precise, we must keep our applications of superadditivity “in the integers”.
The claimed asymptotic for $t=t(m)$ is equivalent to $m=(1+O(t^{-1/2}))(t^{3}/3)^{1/2}$. Set
$M=M(t):=(1+Ct^{-1/2})(t^{3}/3)^{1/2},$
for a universal constant $C>0$ to be specified later; we will establish the claim by showing that $\sum_{i=1}^{t}s(i)<s(M)$. Notice that for $1\leq i\leq t$, superadditivity gives
$x(M)\geq\lfloor M/i\rfloor\cdot x(i)=(1+O(t^{-1/2}))\frac{M}{i}\cdot x(i)$
(using $M/i\gg t^{1/2}$ to bound the error term). Thus (due to the positivity of the $y(i)$’s) we have
$\sum_{i=1}^{t}x(i)y(i)\leq\sum_{i=1}^{t}(1+O(t^{-1/2}))\frac{i}{M}\cdot x(M)y(i)=(1+O(t^{-1/2}))M^{-1}x(M)\sum_{i=1}^{t}iy(i).$
We can write the last sum as
(4) $\underbrace{y(1)}_{1\text{ term}}+\underbrace{y(2)+y(2)}_{2\text{ terms}}+\cdots+\underbrace{y(t)+\cdots+y(t)}_{t\text{ terms}}.$
With this in mind, consider the sequence with $1$ term equal to $1$, then $2$ terms equal to $2$, and so on, up to $t$ terms equal to $t$. Using a greedy algorithm, we partition this sequence into several subsequences each with sum between Mt+1M-t+1 and $M$, and one final subsequence with sum at most $M$; this partition has at most
$\frac{\sum_{i=1}^{t}i^{2}}{M-t+1}+1=(1+O(t^{-1/2}))\frac{t^{3}/3}{M}$
parts. For each part of the partition, the corresponding terms in (4) sum to at most $y(M)$ by superadditivity. In total, we obtain
$\sum_{i=1}^{t}x(i)y(i)\leq(1+O(t^{-1/2}))M^{-1}x(M)\cdot(1+O(t^{-1/2}))\frac{t^{3}/3}{M}\cdot y(M)=\frac{1+O(t^{-1/2})}{(1+Ct^{-1/2})^{2}}x(M)y(M).$
The right-hand side is at most $x(M)y(M)$ as long as $C$ is sufficiently large. This establishes the claim and completes the proof of the lemma. ∎
Applying Lemma 2.2 to the subsequences extracted in Section 2.1 (with $m:=\min(r,r^{\prime})$) completes the proof of Theorem 1.1.
Remark 2.3.
The proof of Lemma 2.2 can be generalized to show that for each $d\geq 2$, there is a constant $\kappa(d)>0$ such that the following holds: For any $d$ superadditive sequences $x_{1}(i),\ldots,x_{d}(i)$ of positive reals, the set of numbers $s(i):=x_{1}(i)x_{2}(i)\cdots x_{d}(i)$ has at least $\kappa(d)m^{d+1}$ distinct subset sums. At this point, it is unclear to us if this has an interpretation in terms of sums of the form
$S(\mathbf{a}_{1},\dots,\mathbf{a}_{d})=\left\{\sum_{i=1}^{n}a_{1,\pi_{1}(i)}a_{2,\pi_{2}(i)}\cdots a_{d,\pi_{d}(i)}:\pi_{1},\pi_{2},\dots,\pi_{d}\in\mathfrak{S}_{n}\right\}.$
where each $\mathbf{a}_{i}=(a_{i,1},\dots,a_{i,n})$ has distinct coordinates. Nevertheless, we conjecture that, generalizing our result when $d=2$, such a set has $\Omega(n^{d+1})$ different permutation-twisted dot products.
3. Acknowledgments
Colin Defant was supported by a Benjamin Peirce Fellowship at Harvard University. Noah Kravitz was supported in part by a NSF Mathematical Sciences Postdoctoral Research Fellowship under grant DMS-2501336. We thank Roger Lidón Ardanuy for posing this question a few years ago, and Daniel Zhu for helpful discussions.

### v2

@@
 We show that this sum always assumes at least $\Omega(n^{3})$ distinct values, which is optimal.
-This “support” bound complements recent work of Do, Nguyen, Phan, Tran, and Vu on the anticoncentration properties of $S$ when $\pi$ is chosen uniformly at random.
+This “support” bound complements recent work of Do, Nguyen, Phan, Tran, and Vu, and of Hunter, Pohoata, and Zhu on the anticoncentration properties of $S$ when $\pi$ is chosen uniformly at random.
 1. Introduction
@@
 which is the “dot product of $\mathbf{a},\mathbf{b}$ twisted by $\pi$”.
-Motivated by Littlewood–Offord theory and a geometric problem of Pawlowski [P24], Do, Nguyen, Phan, Tran, and Vu [DNPTV25] recently raised the problem of studying the anticoncentration properties of the random variable $S(\mathbf{a},\mathbf{b};\pi)$ where $\pi$ is chosen uniformly at random.
+Motivated by Littlewood–Offord theory and a geometric problem of Pawlowski [P24], Do, Nguyen, Phan, Tran, and Vu [DNPTV25], and independently Hunter, Pohoata, and Zhu [HPZ26] recently raised the problem of studying the anticoncentration properties of the random variable $S(\mathbf{a},\mathbf{b};\pi)$ where $\pi$ is chosen uniformly at random.
 One of their main results states that if the $a_{i}$’s are all distinct and the $b_{i}$’s are all distinct, then each possible value of the sum is assumed with probability at most $O(n^{-5/2}\log n)$; this bound is optimal up to the logarithm (which likely can be removed).
-This work has prompted a small flurry of activity [ABDV26, HPZ26].
+The paper [ABDV26] has rounded out this small flurry of recent activity.
 We are interested in the set
 $\mathcal{S}(\mathbf{a},\mathbf{b}):=\left\{S(\mathbf{a},\mathbf{b};\pi):\pi\in\mathfrak{S}_{n}\right\}$
-of all possible sums as $\pi$ ranges; this is of course precisely the support of Do, Nguyen, Phan, Tran, and Vu’s random variable $S(\mathbf{a},\mathbf{b};\pi)$.
+of all possible sums as $\pi$ ranges; this is of course precisely the support of the random variable $S(\mathbf{a},\mathbf{b};\pi)$.
 Our main result is that if the $a_{i}$’s are all distinct and the $b_{i}$’s are all distinct, then there are at least $\Omega(n^{3})$ distinct sums.
@@
 Our proof of Theorem 1.1 is elementary and fairly short.
-As it turns out, we are able to obtain $\Omega(n^{3})$ distinct sums using a somewhat small set of permutations, namely, the permutations obtained as an involution times a fixed “base” permutation.
+As it turns out, we are able to obtain $\Omega(n^{3})$ distinct sums using a somewhat small set of permutations, namely, the permutations obtained as involutions multiplied by a fixed “base” permutation.
 Pohoata [P26] has recently obtained an independent proof of Theorem 1.1 using more complicated methods.
@@
 2.1. Constructing the involutions
-The first step is identifying a family of disjoint pairs $(j_{i},k_{i})$ of elements of $[n]$ such that the corresponding differences $a_{k_{i}}-a_{j_{i}}$ “see” the various scales of $\mathbf{a}$; and likewise for $\mathbf{b}$.
+The first step is identifying a family of disjoint pairs $(j_{i},k_{i})$ of elements of $[n]$ such that the corresponding differences $a_{k_{i}}-a_{j_{i}}$ “see” the various scales of $\mathbf{a}$, and likewise for $\mathbf{b}$.
 We achieve this iteratively, as follows.
@@
 Write $i_{1}<i_{2}<\cdots<i_{u+v-1}$ for the elements of $\mathcal{U}_{u+v}$ between $j_{u+v}$ and $k_{u+v}$.
-All of these elements are also contained in $\mathcal{U}_{u}$ and $\mathcal{U}_{v}$, so the greedy nature of our inductive procedure guarantees that $x(u)\leq a_{i_{u}}-a_{j_{u+v}}$ and $x(v)\leq a_{k_{u+v}}-a_{i_{u}}$, whence $x(u)+x(v)\leq a_{k_{u+v}}-a_{j_{u+v}}=x(u+v)$.
+All of these elements are also in $\mathcal{U}_{u-1}$ and $\mathcal{U}_{v-1}$, so the greedy nature of our inductive procedure guarantees that $x(u)\leq a_{i_{u}}-a_{j_{u+v}}$ and $x(v)\leq a_{k_{u+v}}-a_{i_{u}}$, whence $x(u)+x(v)\leq a_{k_{u+v}}-a_{j_{u+v}}=x(u+v)$.
 ∎